Steven deposits $200 at the end of every month and withdraws $500 at the end of every 6 months. The interest rate is 6% per year, compounded semiannually, and intra-period interest is not paid. Steven wants to know how much money would be in his account at the end of 4 years. What will be the future value
Answer #1

Deposit every month = 200

In six months deposit = 6*200 = 1200

Withdrawl every 6 months =500

Net deposit in six months = 1200-500 = 700

i=6% per year compounded semi annually

i = 0.06/2 =0.03 for half year

A = 700 per period, t=4 yrs = 4*2 = 8 periods

Treating this case as uniform series with 700 payment in every period and i=0.03 per period and total periods =8 and Using Factor (F/A, i%, t)

we get Future value = 700 (F/A, 3%, 8)

=700*8.892336 = 6224.63